OpenStudy (anonymous):

in triangle EFG, m angle E=30 degrees, m angle F= 60 degrees adn m angle G= 90 degrees, which of the following statements are true? please help with explanation! thanks

OpenStudy (campbell_st):

I think you need to post the answer choices

OpenStudy (anonymous):

EG=2*FG EF=2*FG FG=(sqrt3/2)*EF EG=(sqrt3/2)*EF EG=sqrt3 * FG EF= sqrt3 *FG

OpenStudy (anonymous):

sorry was typing them out

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok so its an exact value triangle here is the diagram with side measurements  hope it helps

OpenStudy (anonymous):

oh! ok seeing definitely helps!

OpenStudy (anonymous):

so it's answers 2 3 and 4?

OpenStudy (anonymous):

@campbell_st